1. A coil has inductance of 0.318H and aresistance of 200V .a capacitor of value 3.18Micro farads is connected in parallel with a coil of a sinusoidal supply of variable frequency.calculate the frequency at which the supply of current and voltage are in phase.
1
Expert's answer
2022-03-28T14:25:09-0400

We know that

"X_L=X_c"

"f=\frac{1}{2\pi}\sqrt\frac{1}{LC}"

"f=\frac{1}{2\times3.14}\sqrt\frac{1}{0.318\times3.18\times10^{-6}}"

"f=\frac{994.42}{6.28}=158.34Hz"